If the temperature is decreased to 5 °C, what is the volume of the sample if the pressure is increased to 96.0 kPa?

1
Expert's answer
2022-05-24T05:37:02-0400

In gases, the change in volume is directly proportional to temperature and inversely proportional to pressure. That is,

V / V0 = T / T0 = P0 / P

in the formula:

V – final volume = X

V0 – initial volume = 22,4 l

T – final temperature = 5° C = 278 K

T0 –initial temperature = 20° C = 298 K

P – final pressure = 96 kPa

P0 – initial pressure = 101,325 kPa


V = P0 * V0 * T / P * T0

V = 101,325 * 22,4 * 278 / 96 * 298 = 22,06 l

comment: Since the initial conditions and gas volume were not given in the assignment, 1 mole of gas under normal conditions was taken.

ANSWER: 22,06 litres